# GeniL: A Multilingual Dataset on Generalizing Language

### v1

@@ 3.2 Annotating Generalization in Language
We assigned 3 annotators to label each sentence. Annotators for each language were native speakers, with the exception of English annotations conducted in India, with English proficiency as a selection criteria. For Spanish and Portuguese, we sought annotations from native speakers in Mexico and Brazil, respectively, since we chose source tuples from SGM corresponding to those countries. Annotators were recruited through a proprietary platform, and compensated in accordance to their local law, and were informed of the intended use of their annotations. The annotation process was closely monitored by the authors and recruiters fluent in the respective languages. We ensured task clarity by conducting initial annotation tests (with 50 items) and iteratively enhancing the annotation guide to respond to any recurring ambiguities (see “Notes” paragraph in Appendix A.1). Inter-annotator agreement rates are provided in Table 2, demonstrating moderate to substantial agreement across different languages. For the analyses presented in this paper, annotations were aggregated, and sentences were classified as “Generalizing” if a majority of annotators concurred. However, following (Prabhakaran et al., 2021) we will release individual annotations to enable future studies on subjective differences.
@@ 4.1 Generalization likelihood across associations, identities, and languages
One of the core motivations for building GeniL is that the mere co-occurrence of an identity term and attribute in a sentence doesn’t always convey a generalization context. In fact, our dataset demonstrates that only a very small percentage of sentences with both the identity term and attribute are generalizing in nature: on average 5.9% ($SD$ = 2.4%) of our sentences across all languages are labeled as G, with the highest of 8.7% for Bengali sentences and the lowest of 1.9% for Hindi – see Table 2). In other words, any co-occurrence based approaches to estimate the extent of generalization will be over-estimating by a factor of 10 or more. Another related question is whether the rate at which each (identity term, attribute) pair occurs in generalizing context is similar — in which case one could simply scale down the estimates by a factor of 10. To answer this, we computed the mean and standard deviation for the generalization likelihood of each association as well as each identity term. We observe significant variance in generalization likelihood among different associations (Figure 2). In fact, different identity terms also have huge variance in generalization likelihood suggesting that any co-occurrence based approaches are likely to incur false positives at different rates for different identity terms, which could further introduce undesirable biases. It is also important to note that there is considerable variation in generalization likelihood across different languages. Collectively, these results provide empirical support for the need for assessing stereotypes in the linguistic context of generalization for accurate evaluation.
@@ 4.2 Trends in Stereotypical Generalizations
We now delve deeper into the subset of associations that were validated to be stereotypical in respective societies as per SGM (Bhutani et al., 2024). Here, we treat any association that was labeled as an existing stereotype by at least 2 out of 3 annotators in SGM to be a stereotypical association. The generalization likelihood is higher for associations that are identified as stereotypical in SGM (6.5%) compared to those that are not (4.8%). Similar to the general trend, stereotypical associations are also more likely to appear in non-generalizing language compared to generalizing contexts, across all languages (see Figure 3(a)). We further explore the ratio of offensive attributes being used in generalizing contexts across different languages. Figure 3(b) shows the distribution of offensive attributes across generalizing sentences. Across different languages, when sentences mention an offensive stereotypical association they are on average 7.8% (SD = 3.9%) likely to be generalizing (even higher rate than just all stereotypes). These findings substantiates our claim that that filtering or evaluating text based on co-occurrences of stereotypical association is a biased approach, and attending to the linguistic context in which a stereotypical associations appear has vital importance for detecting stereotypical language and downstream harms.
@@ 5 Multilingual Classifiers for Generalizing Language
  • en+translated: augmenting English data with (machine) translations into the target language. This setting also emulates the case where there is only English GeniL data, but mitigating the multilingual gap through automated translations.
@@ 6.1 Limitations
In this paper, we focus on cases of explicit generalizations that happen in the same sentence. However, generalization may sometimes happen over multiple sentences. For instances the two sentences “I met a French man at the train station. Of course, he was rude!” are together perpetuating the stereotype that French people are rude, however this does not happen in the same sentence. GeniL will miss such instances, despite them being made explicit in language. Future work should expand into such discourse level generalizations. We also do not capture implicit generalizations that do not have both the identity term and the attribute mentioned in text. Finally, whether or not a sentence has generalizing language has some level of subjectivity. While we ensured gender balance in our rater pool, it may have skews along other socio-demographic axes. We used majority vote in this work to arrive at a ground truth label, however future work should look into individual and group level differences. We release individual annotations to enable any such follow up studies. While the main contribution of this paper is the task and the dataset, our classifier experiments suggest that the task of detecting generalization as well as distinguishing the two kinds are both non-trivial tasks computationally. The overall PR-AUC is 58.7 which suggest ample opportunities for improvement. In particular, the performance is the poorest for Bengali, followed by Indonesian and Arabic (all three posted PR-AUC less than 50), which are also the languages that obtained the lowest inter-rater agreement. Future work should look into improving both the data quality and model performance in these languages, as well as expand coverage to other low resource languages. Future work should also expand GeniL to other languages with diverse morpho-syntactic features to ensure more robust coverage across language families.
@@
We build GeniL dataset and classifiers with the intention to be used for evaluation purposes. By design, the dataset contains a high incidence of potentially problematic generalizations about groups of people. While this corpus enables large-scale and finer-grained stereotype evaluation, the concentration of such sentences with potentially harmful associations may be triggering for certain individuals. Hence care must (and will) be taken in how the data is disseminated and used downstream. There is a risk of misuse wherein the GeniL tagger not detecting an instance of generalization is deemed as an absence of generalization. In addition to the error in recall, it is also important to remember that GeniL focuses only on explicit generalizations, and instances of implict generalization may still be present. Finally, there is a potential for malicious use where someone could use the GeniL dataset to intentionally generate generalizing language targeting specific identity groups.
@@ Acknowledgements
We thank Akshita Jha for early discussions on this project. We thank Kevin Robinson and Kathy Meier-Hellstern for useful feedback on early versions of this manuscript.
@@ A.1 Annotation Manual
Sentence
Explanation
Label
1
Poverty spreads from cities to countryside despite the people of Laos’ hardworking perseverance.
The sentence states that all people of Laos are hardworking.
PG
2
It turns out the French aren’t just good at making wine, but they can make a mean cider as well!
The sentence argues for a generalization (French are good at making wine and cider)
PG
3
For those who think the Irish are a hard-drinking but jolly race, this play will be an eye-opener.
The sentence suggests that a generalization exists (Irish are har-drinking and jolly), but does not promote it.
MG
4
We all enjoyed an amazing French Day last week with lots of singing as we learnt colours and numbers in French and ate cheese and bread in our lovely blue, white and red outfits!
NG
5
Travelers will find Turks to be exceptionally gracious hosts.
The sentence implies that Turks are gracious hosts.
PG
6
Representative Maritza Davila said that the Turks are the most diligent among the ethnic communities in New York.
The sentence mentions a generalization made by a person, and does not promote it.
MG
7
If you want to add some new moves to your wrestling arsenal then be sure to pay a visit to Randall Lovikiv, the hard drinking Russian.
The sentence mentions a hard drinking Russian, does not generalize any attribute to all Russians.
NG
@@ A.2 Distribution of Generalizing language
Table 5 shows the number of GeniL sentences across different languages that mention a stereotypical or non-stereotypical association (based on Multilingial SeeGULL Bhutani et al. (2024)), separated by their generalizing (G) or non-generalizing (NG) context, determined in our annotations.

### v2

@@ 3.2 Annotating Generalization in Language
We assigned 3 annotators to label each sentence. Annotators for each language were native speakers, with the exception of English annotations conducted in India, with English proficiency as a selection criteria. For Spanish, Portuguese, and Arabic, we sought annotations from native speakers in Mexico, Brazil, and Saudi Arabia, respectively, since we chose source tuples from SGM corresponding to those countries. Annotators were recruited through a proprietary platform, and compensated in accordance to their local law, and were informed of the intended use of their annotations. The annotation process was closely monitored by the authors and recruiters fluent in the respective languages. We ensured task clarity by conducting initial annotation tests (with 50 items) and iteratively enhancing the annotation guide to respond to any recurring ambiguities (see “Notes” paragraph in Appendix A.1). Inter-annotator agreement rates are provided in Table 2, demonstrating moderate to substantial agreement across different languages. For the analyses presented in this paper, annotations were aggregated, and sentences were classified as “Generalizing” if a majority of annotators concurred. However, following (Prabhakaran et al., 2021) we will release individual annotations to enable future studies on subjective differences.
@@ 4.1 Generalization likelihood across associations, identities, and languages
One of the core motivations for building GeniL is that the mere co-occurrence of an identity term and attribute in a sentence does not always convey a generalization context. In fact, our dataset demonstrates that only a very small percentage of sentences with both the identity term and attribute are generalizing in nature: on average 5.9% ($SD$ = 2.4%) of our sentences across all languages are labeled as G, with the highest of 8.7% for Bengali sentences and the lowest of 1.9% for Hindi – see Table 2). In other words, any co-occurrence based approaches to estimate the extent of generalization will be over-estimating by a factor of 10 or more. Another related question is whether the rate at which each (identity term, attribute) pair occurs in generalizing context is similar — in which case one could simply scale down the estimates by a factor of 10. To answer this, we computed the mean and standard deviation for the generalization likelihood of each association as well as each identity term. We observe significant variance in generalization likelihood among different associations (Figure 2, 5). In fact, different identity terms also have huge variance in generalization likelihood suggesting that any co-occurrence based approaches are likely to incur false positives at different rates for different identity terms, which could further introduce undesirable biases. It is also important to note that there is considerable variation in generalization likelihood across different languages. Collectively, these results provide empirical support for the need for assessing stereotypes in the linguistic context of generalization for accurate evaluation.
@@ 4.2 Trends in stereotypical generalizations
We now delve deeper into the subset of associations that were validated to be stereotypical in respective regions as per SGM (Bhutani et al., 2024). Here, we treat any association that was labeled as an existing stereotype by at least 2 out of 3 annotators in SGM to be a stereotypical association. The generalization likelihood is higher for associations that are identified as stereotypical in SGM (6.5%) compared to those that are not (4.8%). Similar to the general trend, stereotypical associations are also more likely to appear in non-generalizing language compared to generalizing contexts, across all languages (see Figure 3(a)). The varying distributions observed in Figure 3 can be interpreted as potential evidence of the various manifestation of stereotypes and generalizations in different linguistic and cultural contexts. It is possible that generalizations are more prevalent, explicit, or easily detectable in French, Bengali, or Malay sources compared to others. Alternatively, the observed variations might reflect variations in the representation of each language in mC4, including factors such as sample size, or source selection. We further explore the ratio of offensive attributes being used in generalizing contexts across different languages. Figure 3(b) shows the distribution of offensive attributes across generalizing sentences. Across different languages, when sentences mention an offensive stereotypical association they are on average 7.8% (SD = 3.9%) likely to be generalizing (even higher rate than just all stereotypes). These findings substantiates our claim that filtering or evaluating text based on co-occurrences of stereotypical association is a biased approach, and attending to the linguistic context, in which a stereotypical association appears, has vital importance for detecting stereotypical language and downstream harms.
@@ 5 Multilingual Classifiers for Generalizing Language
  • en+translated: augmenting English data with (machine333using the publicly available Google Translate API) translations into the target language. This setting also emulates the case where there is only English GeniL data, but mitigating the multilingual gap through automated translations.
@@ 6.1 Limitations
In this paper, we focus on cases of explicit generalizations that happen in the same sentence. However, generalization may sometimes happen over multiple sentences. For instances the two sentences “I met a French man at the train station. Of course, he was rude!” are together perpetuating the stereotype that French people are rude, however this does not happen in the same sentence. GeniL will miss such instances, despite them being made explicit in language. Future work should expand into such discourse level generalizations. We also do not capture implicit generalizations that do not have both the identity term and the attribute mentioned in text. GeniL is not intended to be a full representation of all potentially stereotypical or generalizing sentences in the target languages. It is a focused lens on the diverse linguistic contexts of generalization. While SGM, a large validated dataset of stereotypes, is an excellent starting point, we do recognize its inherent limitations and biases. Moreover, whether or not a sentence has generalizing language has some level of subjectivity. While we ensured gender balance in our rater pool, it may have skews along other socio-demographic axes. We used majority vote in this work to arrive at a ground truth label, however future work should look into individual and group level differences. We release individual annotations to enable any such follow up studies. While the main contribution of this paper is the task and the dataset, our classifier experiments suggest that the task of detecting generalization as well as distinguishing the two kinds are both non-trivial tasks computationally. The overall PR-AUC is 58.7 which suggest ample opportunities for improvement. In particular, the performance is the poorest for Bengali, followed by Indonesian and Arabic (all three posted PR-AUC less than 50), which are also the languages that obtained the lowest inter-rater agreement. It can be interpreted that this task is more challenging for humans to annotate in certain languages, due to factors such as linguistic nuances or cultural differences. This in turn could make it more difficult for the model effectively in those languages. Future work should look into improving both the data quality and model performance in these languages, as well as expand coverage to other low resource languages. Future work should also expand GeniL to other languages with diverse morpho-syntactic features to ensure more robust coverage across language families.
@@
We build GeniL dataset and classifiers with the intention to be used for evaluation purposes. By design, the dataset contains a high incidence of potentially problematic generalizations about groups of people. While this corpus enables large-scale and finer-grained stereotype evaluation, the concentration of such sentences with potentially harmful associations may be triggering for certain individuals. Hence care must (and will) be taken in how the data is disseminated and used downstream. There is a risk of misuse wherein the GeniL tagger not detecting an instance of generalization is deemed as an absence of generalization. In addition to the error in recall, it is also important to remember that GeniL focuses only on explicit generalizations, and instances of implicit generalization may still be present. Finally, there is a potential for malicious use where someone could use the GeniL dataset to intentionally generate generalizing language targeting specific identity groups.
@@ Acknowledgements
We thank Akshita Jha for early discussions on this project. We also appreciate Kevin Robinson and Kathy Meier-Hellstern for useful feedback on early versions of this manuscript. We thank the annotation team for facilitating our data work.
@@ A.1 Annotation Manual
Sentence Explanation Label
1 Poverty spreads from cities to countryside despite the people of Laos’ hardworking perseverance. The sentence states that all people of Laos are hardworking. PG
2 It turns out the French aren’t just good at making wine, but they can make a mean cider as well! The sentence argues for a generalization (French are good at making wine and cider) PG
3 For those who think the Irish are a hard-drinking but jolly race, this play will be an eye-opener. The sentence suggests that a generalization exists (Irish are har-drinking and jolly), but does not promote it. MG
4 We all enjoyed an amazing French Day last week with lots of singing as we learnt colours and numbers in French and ate cheese and bread in our lovely blue, white and red outfits! NG
5 Travelers will find Turks to be exceptionally gracious hosts. The sentence implies that Turks are gracious hosts. PG
6 Representative Maritza Davila said that the Turks are the most diligent among the ethnic communities in New York. The sentence mentions a generalization made by a person, and does not promote it. MG
7 If you want to add some new moves to your wrestling arsenal then be sure to pay a visit to Randall Lovikiv, the hard drinking Russian. The sentence mentions a hard drinking Russian, does not generalize any attribute to all Russians. NG
@@
Figure 5: The frequency with which associations and identities appear in generalizing contexts varies across languages. This is the same info as Figure 2 using a box plot
@@ A.2 Distribution of Generalizing language
Table 5 shows the number of GeniL sentences across different languages that mention a stereotypical or non-stereotypical association (SGM), separated by their generalizing (G) or non-generalizing (NG) context, determined in our annotations.